Trisha, who is single and has taxable income of $300,000, sells the following capital assets in 2019 with gains and losses as shown. With accurate calculations please.

Asset Gain or (Loss) Holding Period A $15,000 15 months B 7,000 20 months C (3,000) 14 months

a. Determine Trisha’s increase in income tax liability as a result of the three sales. All assets are stock held for investment. Ignore the effect of increasing AGI on deductions and phaseout amounts.

b. Determine her increase in income tax liability if the holding period for asset B is 8 months.

c. Determine her increase in income tax liability if the holding periods are the same as in the Part a but asset B is an antique clock.

d. Determine her increase in income tax liability if her taxable income is $700,000.
Answer #1

Trisha, who is single and has taxable income of $300,000, sells the following capital assets in 2019 with gains and losses as shown. With accurate calculations please.

Question :- Trisha , who is single and has taxable income of $ 300.000. Sells the following capital assets in 2019 with gainsTotal Gain = $ 15000+ 7000 - 3000 = $19000. Taxable amount = $ 19000* 0.15 - $ 2850. (b. Determine hoor incorease in income f(C) Determine her increase in income tax liability if the holding Periods are the same as in the part a but asset B is an antTotal Gain = $15.000 + 7000 - 3000 = $19000. Taxable amount = $19000 * 0.20 = $ 2800.